0612024697:

A company's monthly profit, P, from a product is given by P=−x2+105x−1050, where x is the price of the product in dollars. What is the lowest price of the product, in dollars, that gives a monthly profit of $1550?

Florisalreadytaken:

just another quadratic formula \[ 1550=−x^2+105x−1050 \] can you work that out?

0612024697:

how do you work out 2,600=-x squared +105x

Florisalreadytaken:

how about: \[ 0=x^2-105x+2600 \]

0612024697:

I added 1050 to both sides and got 2,600=-x squared +105x

Florisalreadytaken:

yeah nah you want to equalize it by 0, so that you can then solve it ok so we can see that \( a=1 ~~~ ; ~~~ b=-105 ~~~ ;~~~ c=2600 \) plug that information into \( \frac{-b\pm\sqrt{b^2-4ac}}{2a} \)

0612024697:

ok

0612024697:

Thank you

0612024697:

I got 65 and 40

0612024697:

I got it right

Florisalreadytaken:

youre right -- i will just do the latex cuz it took me a while and i dont want to delete it \[ x= \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[ x= \frac{-(-105)\pm\sqrt{(-105)^2-4\cdot 1 \cdot 2600}}{2\cdot 1 } \] \[x= \frac{-(-105)\pm\sqrt{(-105)^2-10400}}{2\cdot 1 } \] \[ x= \frac{-(-105)\pm\sqrt{11025-10400}}{2\cdot 1 } \] \[x= \frac{-(-105)\pm\sqrt{625}}{2\cdot 1 } \] \[ x= \frac{-(-105)\pm\sqrt{625}}{2\cdot 1 } \] \[ x=\frac{-(-105)\pm\sqrt{25^2}}{2\cdot 1 } \] \[ x= \frac{105\pm25}{2\cdot 1 } \] \[ x_1=65 ~~~~~~~~~~ x_2=40 \] now back to the question "What is the `lowest price` of the product, in dollars, that gives a monthly profit of $1550? " so the answer would be what?

0612024697:

40$

Florisalreadytaken:

there you go!